SUR operational status:
The station operator replaced the SUR00 Q330 digitizer on day 2020,225 (Aug.12) and the GPS clock has locked.
There was a timing issue with the GPS clock from the SUR00 Q330 digitizer. While the clock quality recorded was in normal operating range, there was an offset and drift in timing from mid January to mid May. A reboot of the Q330 digitizer enabled the GPS to re-lock and correct timing was restored.
There was high background noise on the horizontal channels from the STS-1 sensors and episensor channels starting day 2017,115. Data quality improved on day 2017,129 after the episensor was disconnected. IDA engineers are working with the station operator to determine if the episensor needs to be replaced to resolve this issue.
Data acquisition resumed when power was restored.
IDA station SUR has been upgraded to a Quanterra Q330HR system. Seismometers include an STS-1, STS-2, and an episensor strong motion detector.
The wdo channel from the paroscientific microbarometer was not being acquired from about day 2008,186 to 2008,204 due to communication problems between the microbarometer and the digi.
The wdo channel from the paroscientific microbarometer was not being acquired from about day 2008,141 to 2008,149 following a power outage at the station. The channel was restored after the digi was reset.
The East-West channels from the STS-1 seismometer were nonseismic from about day 2008,048-05:03 to 2008,051-19:31. The channels were restored after a reset of the DAS.
The station is back online as of about day 2007,345 after the station operator cycled power to the GPS clock.
The station is down as of about day 2007,341 due to problems with the GPS clock that are yet to be determined.
The station is back online as of about day 2007,240-12:05 after the station operator checked the power system and cycled power to the DAS.
The station is down as of about day 235-03:21 due to problems with the DAS that are yet to be determined.
The wdo channel is restored as of about day 2007,226-14:29 after the station operator replaced the microbarograph.
The East-West channels from the STS-1 seismometer have been restored as of about day 165-08:50 after an IDA engineer visited the station and replaced the circuit board in the DAS.
The station was down from day 164-14:41 to 166-07:32 due to problems with the communication link between the digitizer and the ISP. The station operator reconnected the link and data acquisition resumed.
The station operator restored the vertical channels for the STS-1 by replacing the faulty circuit board with the functional board from the E-W channels. The wdo channel is out.
The station is up after the station operator replaced the DAS. The vertical channels for the STS-1 seismometer are out due to a faulty circuit board in the DAS. The station operator has been requested to remove the faulty board and insert one of the functional boards from the horizontal channels’ slots into the vertical channel’s slot.
The station is down as of day 334 (Nov.30) due to communication problems between the DAS and the ISP.
The BBN channels for the STS-1 sensor are out as of 2005:244-06:13 due to a faulty circuit board in the DAS. The station operator has been instructed and has removed the faulty circuit board. Currently the BBN and BBE channels are out.
Station was down day 193-18:26 to 196-07:38 due to problems with communication between the DAS and ISP. The station came back up after the station operator cycled power to the DAS.
Station is up again after the station operator power cycled the digitizer.
DAS stops sending data. Station operator has been notified.
The data acquisition system has been upgraded to a MK7-ISP system. An IDA engineer installed a Paroscientific microbarograph and replaced the main and vault power supplies, GPS receiver, and station processor. A faulty RT303 card was removed from the DAS for the BBZ channels and was replaced with an operational RT303 card from the BBE channels.
Station is back up after the station operator removed the faulty RT303 board for the digitizer’s BBN channels.
Station is down due to problems with the DAS/ARS. There is a faulty RT303 board in the digitizer.
There are several gaps in recent NRTS data due to communication errors with the Auxiliary Recording System (ARS). After the ARS was rebooted at 2003:331-17:02, (11/27) continuous NRTS data to the disk loop resumed. However, there are timing problems with the data due to a faulty RT303 board for the digitizer’s BBN channels.
Telemetry at the station was restored after the operator installed a replacement workstation shipped to the site from UCSD.
Data streams from the STS2 sensor returned to normal after the DAS boards were reseated by the operator. When the EPROMs were replaced on 7/15/02, the boards had been improperly re-inserted into the DAS, which resulted in loss of all STS2 and STS1 mass position data.
Data streams from the STS2 sensor are non-seismic.
Station operating normally.